### v2

# Number of local minima in discrete-time fractional Brownian motion

Abstract
The analysis of local minima in time series data and random landscapes is essential across numerous scientific disciplines, offering critical insights into system dynamics. Recently, Kundu, Majumdar, and Schehr derived the exact distribution of the number of local minima for a broad class of Markovian symmetric walks [Phys. Rev. E 110, 024137 (2024)]; however, many real-world systems are non-Markovian, typically due to interactions with possibly hidden degrees of freedom. This work investigates the statistical properties of local minima in discrete-time samples of fractional Brownian motion (fBm), a non-Markovian Gaussian process with stationary increments, widely used to model complex, anomalous diffusion phenomena. We derive a complete asymptotic characterization of the fluctuations of the number of local minima $m_{N}$ in an $N$-step discrete-time fBm. We show that the fluctuations of $m_{N}$ exhibit a sharp transition at the Hurst exponent $H=3/4$: for $H\leq 3/4$ they satisfy a central limit theorem with Gaussian limiting law, whereas for $H>3/4$ they converge to a non-Gaussian Rosenblatt process. The convergence at the process level gives us full statistical description at all times. We exemplify it on the covariance of the rescaled minima process, which displays two qualitatively distinct regimes matching Brownian and Rosenblatt covariances on either side of this threshold. Our analysis relies on a Hermite/Wick decomposition of the local-minimum indicator, which isolates a quadratic functional of an effective long-memory mode as the unique driver of the anomalous statistics. These results identify the count of local minima as a simple and robust diagnostic of long-range dependence in non-Markovian Gaussian processes, a conclusion supported by numerical simulations.
The identification and analysis of local minima in time series and random landscapes play a fundamental role across various scientific fields [1, 2]. Their occurrence and distribution often reveal crucial insights about the system, whether biological, environmental, financial, or physical. In biomedical applications, local minima in electrocardiogram signals are used to detect cardiac disorders [3] and to identify sleep phases in neuroimaging [4]. Similarly, climate scientists aim to predict minimum daily temperatures for frost forecasting [5], financial analysts examine local extrema to determine market trends [6, 7], and in signal processing, local minima are used to decompose complex signals [8]; finally, the analysis of local minima finds a wide range of applications in physics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] — for instance, the nummber of local minima in energy landscapes are critical to understanding phase transitions [11, 12, 13, 9, 14, 15, 10], as their abundance of such minima significantly influences the low-temperature behavior of spin glasses [18, 19].
Figure 1: fBm trajectories of Hurst exponents $H=1/4$ and $3/4$ (blue line in top and bottom figures, respectively), sampled at integer-valued times (red line). Black dots indicate the local minima in the sampled trajectories, the number of which strongly depends on the Hurst exponent.
Recently, Kundu, Majumdar, and Schehr derived the exact distribution of the number of local minima for a broad class of Markovian symmetric walks [21, 22], which is expressed explicitly in terms of combinatorial coefficients and converges to a Gaussian distribution with mean $N/4$ and variance $N/16$ as the number of steps $N$ grows. However, many real-world systems are non-Markovian, typically due to interactions with possibly hidden degrees of freedom [23, 24, 25, 26]. For example, in the context of the above-mentioned applications, heartbeat and electroencephalogram signals [27, 28] as well as pricing options [29, 30], are known to exhibit memory. Describing such non-Markovian dynamics is typically a difficult task.
One of the most fundamental approaches to non-Markovian dynamics is fractional Brownian motion (fBm), a symmetric Gaussian process with stationary increments that has long-range correlations [31]. While, strictly speaking, fBm is a continuous, non-smooth process, in real-world applications it is sampled in discrete time. The resulting discrete-time process $X_{0},\dots,X_{N}$ (see Fig. 1) has increments $\phi_{i}=X_{i+1}-X_{i}$, called fractional Gaussian noise (fGn), which follow a multivariate Gaussian distribution:
$P_{N}(\Phi)$ $=\frac{1}{\sqrt{(2\pi)^{N}\det{\bf R}}}\exp\left(-\frac{1}{2}\Phi^{T}{\bf R}^{-1}\Phi\right),$ (1)
where $\Phi=(\phi_{0},\dots,\phi_{N-1})$}, and ${\bf R}=(\rho_{ij})$} is the correlation matrix, whose elements depend only on the distnance between increments, $\rho_{ij}=\rho_{0|i-j|}$ and
$\rho_{0k}=\frac{1}{2}\left(|k-1|^{2H}+|k+1|^{2H}-2k^{2H}\right),$ (2)
where $H$ ($0<H<1$) is the Hurst exponent  [^1] It was shown in Ref. [74] that the discrete-time fBm $X_{N}$ converges to the continuous fBm that is widely used for construction of fBm [75, 76, 77, 78, 53, 79].. The large $N$ behavior of the corresponding variance $\text{Var}(X_{N})$ of the position $X_{N}$ is given by $\text{Var}(X_{N})\propto N^{2H}$, so that, for $H<1/2$, fBm exhibits subdiffusive motion; for H>1/2{H>1/2}, it is superdiffusive; and only for $H=1/2$ is the motion diffusive and Markovian, with $\rho_{0k}=0$.
fBm has broad applications across various fields. It has been shown to effectively model the subdiffusive behavior of telomeres within the cell nucleus [33, 34, 25], as well as the constrained motion of chromosomal loci [35, 36]. The same framework has also proven useful for describing the intracellular transport of lipid granules during early mitosis [37], and for characterizing the dynamics of beads suspended in viscoelastic media [24, 38, 25, 26]. Furthermore, fBm provides insight into the motion of tracer particles in crowded fluidic environments [39], and notably captures the superdiffusive transport of vacuoles in amoeboid cells [40, 25].
In the context of the minima number — which falls within the field of extreme-value statistics — several results have been obtained for fBm, including: survival probability [41, 42, 43, 44, 45], mean first-passage time in confinement [46], time of maximum [47], the statistics of record ages [48], the splitting probability to reach a remote rather than a nearby target [49, 50, 26], and the exploration dynamics of $d$-dimensional spaces [51]. Here we focus on the full characterisation of the minima number $m_{N}$ in fBm, which formally can be written as
$m_{N}$ $=\sum_{i=1}^{N-1}\Theta(-\phi_{i-1})\Theta(\phi_{i}).$ (3)
Here $\Theta(\phi)$ is the Heaviside function, so that the random variable $\Theta(-\phi_{i-1})\Theta(\phi_{i})$ represents change in the slope of $X_{i}$.
The mean of the number of local minima in fBm, studied in [52, 53],
$\langle m_{N}\rangle=\frac{N-1}{4}\left[1-\frac{2}{\pi}\arcsin\rho_{01}\right].$ (4)
is linear for large $N$ as in Markovian case [21]. Meanwhile its prefactor is modified, but it depends only on the nearest-neighbor increment correlation $\rho_{01}$. In fact, it is the same as for the so-called autoregressive model AR(1) [54, 55], which is a discretized Ornstein-Uhlenbeck process [56], and whose correlation matrix is given by:
[…]
Therefore, although the mean $\langle m_{N}\rangle$ deviates from the Markovian behavior (which is obtained by setting $\rho_{01}=0$}), it does not incorporate the key feature of the fBm process — the long-range correlations described by Eq. (2).
Conversely, the variance of the number of local minima, $\text{Var}(m_{N})$, for fBm exhibits a remarkably rich behavior that accounts for all increment correlations $\{\rho_{0k}\}$. Depending on the value of the Hurst exponent $H$,
$\text{Var}(m_{N})\sim$ $\begin{cases}c_{H}N,&H<\frac{3}{4},\\ \frac{9(\sqrt{2}-1)}{64\pi^{2}}N\log N+b_{3/4}N,&H=\frac{3}{4},\\ \frac{4^{1-H}-1}{2\pi^{2}}\frac{H^{2}(2H-1)}{(4H-3)}N^{4H-2},&H>\frac{3}{4}.\end{cases}$ (6)
at leading order in $N$ (see Supplemental Material (SM)  [^2] See Supplemental Material at [url] for detailed calculations and numerical checks of asymptotics for detailed calculations and numerical checks of asymptotics, as well as for definitions of coefficients $c_{H}$ and $b_{3/4}$). We note that the asymptotic scaling exponents of Eq. (6) also follow from Ref. [58], which considers the variance of the number of zero crossings of fGn in a unit interval. Because the fGn $\phi_{i}$ is a discrete derivative of the dicrete-time fBm $X_{i}$, $\phi_{i}=X_{i+1}-X_{i}$, the number of local minima $m_{N}$ of fBm is equal to the number of zero crossings of fGn in a unit interval multiplied by $N/2$ and their variances are related by the factor $N^{2}/4$, respectively. However, in the regime $H>3/4$, which is a key analytic result, Ref. [58] contains an error in the prefactor; our result (6) corrects this prefactor (see SM for two independent proofs). Importantly, this behavior comes from the long-range increment correlations $\rho_{0k}$. Given that $\rho_{0k}\propto k^{2H-2}$ for $k\rightarrow\infty$ and that the correlation function $\rho_{0k}$ is squared in $\text{Var}(m_{N})\propto\sum_{k=2}^{N-2}(N-k-1)\rho_{0k}^{2}$, the correlations become non-summable when 2(2H2)=12(2H-2)=1, i.e. at $H=3/4$, which is not the usual threshold value $H=1/2$ that separates fBm sub- and superdiffusion  [^3] On the level of a path-integral representation, the fBm action involves a kernel with decay exponent 22H2-2H [80]; here the correlations of the minima involve the squared correlations of fBm increments, which are expected to be controlled by an effective kernel with exponent 44H4-4H (which becomes marginal, equal to $1$, at $H=3/4$)..
In the regime $H\leq 3/4$, our work provides further information on $\mathrm{Var}(m_{N})$ beyond that which can be obtained from Ref. [58]: For $H<3/4$, the prefactor $c_{H}$ does not seem to admit a compact analytic expression, as also stated in [58]. Nevertheless, we derive in the SM its behavior close to $H=1/2$ by treating $\epsilon=H-1/2$ as a small parameter, and obtain $c_{H}=\frac{1}{16}+\chi_{1}\epsilon+\chi_{2}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$, with $\chi_{1}=(3\ln 3-4\ln 2)/(4\pi)\approx 0.042$ and $\chi_{2}\approx 0.069$. This shows that non-Markovian effects already enter $\mathrm{Var}(m_{N})$ at first order in a perturbative expansion around Brownian motion. In the marginal case $H=3/4$, we further compute the prefactor of the subleading term and find $b_{3/4}\approx 0.0630$, which is more than ten times larger than the prefactor of the leading term ($\approx 0.00590$). As a result, the crossover to the asymptotic $N\log N$ behavior is extremely slow in practice (see SM).
The strong correlations of $\{\phi_{i}\}$ leading to the anomalously large (superlinear) fluctuations of $\text{Var}(m_{N})$ have a fundamental consequence for the distribution of $m_{N}$, for which the CLT breaks down. We show below this breakdown analytically and confirm it numerically through simulations (see Fig. 2(a))  [^4] The breakdown was also found in the continuous setting of the counterpart zero-crossing problem for fBm studied by Slud, who showed that the CLT fails for $H>3/4$ by representing the number of zero crossings as a second-order Wiener–Itô integral [72, 73].. In the analytic approach, we use the methods Ref. [61] for functionals of Gaussian correlated variables, which allow us to go beyond one-time distribution and to show the convergence of $m_{N}$ to the processes belonging to two different universality classes (Rosenblatt and Brownian motion), separated through the threshold value $H=3/4$.
First step consists in the expansion of $m_{N}$ in the probabilists’ Hermite polynomials $\mathrm{He}_{n}(\phi_{i})$. Using
$\Theta(\pm\phi_{i})=\frac{1}{2}\pm\sum_{\begin{subarray}{c}n\geq 1\\ n\ \mathrm{odd}\end{subarray}}a_{n}\,\mathrm{He}_{n}(\phi_{i}),$ (7)
with $a_{2k+1}=(-1)^{k}/(\sqrt{2\pi}\,(2k+1)\,2^{k}\,k!)$ and $a_{2k}=0$, yields the explicit expansion of $M_{i}\equiv\Theta(-\phi_{i-1})\,\Theta(\phi_{i})$,
$M_{i}$ $=\frac{1}{4}+\sum_{\text{odd }n}\frac{a_{n}}{2}\big(\mathrm{He}_{n}(\phi_{i})-\mathrm{He}_{n}(\phi_{i-1})\big)$ (8)
$-\sum_{\text{odd }m,n}a_{m}a_{n}\,\mathrm{He}_{m}(\phi_{i-1})\,\mathrm{He}_{n}(\phi_{i})$ (9)
The Hermite processes are classified according to the minimal nonzero order in their algebraic expansion in the Hermite polynomials [61]. Equation (8) contains the linear terms proportional to $\mathrm{He}_{1}(\phi_{i})-\mathrm{He}_{1}(\phi_{i-1})$ of rank $1$. However, in the number of minima $m_{N}$, Eq. (3), this contribution is a telescopic difference, leaving only the boundary increments $\phi_{0}$ and $\phi_{N-1}$ and thus does not contribute to $m_{N}$ asymptotically as $N\to\infty$. The term of Eq. (9) having minimal degree, with non-cancelled in $m_{N}$ increments $\{\phi_{i}\}$ for all $i$, is proportional to $\mathrm{He}_{1}(\phi_{i-1})\,\mathrm{He}_{1}(\phi_{i})$. This bilinear form can be expressed in terms of $\mathrm{He}_{2}(\dots)$, as we also show below. Therefore, the $m_{N}$ of Eq. (3) has asymptotically ($N\to\infty$) the Hermite rank $2$. In the regime of strong long-range correlations ($H>3/4$) leading to superlinear $\text{Var}(m_{N})$, these quadratic terms will dominate all fluctuations (here we use Theorem 5.3.3 of Ref. [61] for general functionals of Gaussian correlated variables to show that the higher order terms become irrelevant). In the opposite regime ($H<3/4$) all orders starting from order $2$ matter, but despite of the long-range correlations, the Breuer-Major theorem [62] for functionals of Gaussian correlated variables guarantees that the CLT holds.
[…]
Figure 2: (a) Probability density function (PDF) of centered and normalized $m_{N}$ for N=1024{N=1024}, obtained from simulations (symbols), compared with Gaussian PDF (black line) and (b) for $N=10^{6}$} and various $H>3/4$ compared with the PDF of the negative Rosenblatt variable (lines).
Because the rank $2$ terms drive the CLT breaking, we are going now to analyse Eq. (9), by spanning the quadratic part $M_{i}^{\circ}\equiv M_{i}-\langle M_{i}\rangle$ on the orthogonal vectors of the complete quadratic Wick basis
$B_{1}\equiv\mathrm{He}_{2}(\phi_{i-1})+\mathrm{He}_{2}(\phi_{i})=\phi_{i-1}^{2}+\phi_{i}^{2}-2,$
$B_{2}\equiv\mathrm{He}_{1}(\phi_{i-1})\,\mathrm{He}_{1}(\phi_{i})-\rho=\phi_{i-1}\phi_{i}-\rho,$ (10)
where $\rho=\rho_{01}=\mathrm{Corr}(\phi_{i-1},\phi_{i})=2^{2H-1}-1$. The Wick basis, Eq. (10), has the advantage that any centered quadratic functional of $(\phi_{i-1},\phi_{i})$ can be written as a linear combination of $B_{1}$ and $B_{2}$, e.g.,
$M_{i}^{\circ}=Q_{i}+R_{i},\qquad Q_{i}=c_{1}B_{1}+c_{2}B_{2},$ (11)
where the coefficients $c_{1},c_{2}$ are determined by the orthogonality conditions
$\langle R_{i},B_{j}\rangle=0,\qquad j=1,2,$ (12)
with $\langle A,B\rangle=\mathbb{E}[AB]$. In other words, $Q_{i}$ is the orthogonal projection of $M_{i}^{\circ}$ onto the second order stochastic terms and the remainder $R_{i}$ collects all the components of order $q\neq 2$ (including the rank–1 term that telescopes and all higher orders $q\geq 3$). Solving the projection problem of Eq. (12) yields (see SM for details):
$Q_{i}=\frac{\rho(\phi_{i-1}^{2}+\phi_{i}^{2}-2)-2(\phi_{i-1}\phi_{i}-\rho)}{4\pi\sqrt{1-\rho^{2}}}.$ (13)
[…]
Figure 3: The covariance of the number of minima $\mathbb{E}[Z_{K}(t)Z_{K}(s)]$ (symbols) compared with theory (solid lines), For (a)-(c) the theoretical covariance is $\min(t,s)$ of the Brownian motion and for (d)-(f) is given by Eq. (19) for the Rosenblatt process.
Introducing the normalized sum and difference modes
$V_{i}=\frac{\phi_{i}+\phi_{i-1}}{\sqrt{2(1+\rho)}},\qquad U_{i}=\frac{\phi_{i}-\phi_{i-1}}{\sqrt{2(1-\rho)}},$ (14)
which are standard and independent (for fixed $i$), we can rewrite the quadratic component as
$Q_{i}=-\frac{\sqrt{1-\rho^{2}}}{4\pi}\left[\mathrm{He}_{2}(V_{i})-\mathrm{He}_{2}(U_{i})\right].$ (15)
Summing over $i$, we obtain the fluctuation decomposition
$m_{N}-\langle m_{N}\rangle=-\frac{\sqrt{1-\rho^{2}}}{4\pi}\sum_{i=1}^{N-1}\left[\mathrm{He}_{2}(V_{i})-\mathrm{He}_{2}(U_{i})\right]+\sum_{i=1}^{N-1}R_{i}.$ (16)
The sum of the remainder $R_{i}$ contains (i) boundary increments that do not contribute asymptotically to $m_{N}-\langle m_{N}\rangle$} as $N\to\infty$ at any $H$ and (ii) the fluctuations of higher than quadratic order that are irrelevant for $H>3/4$ (Theorem 5.3.3 of Ref. [61]).
We now focus on the regime (ii), in which the high-$i$ terms linear in $\mathrm{He}_{2}(\dots)$ dominate the sum in Eq. (16), leading to the superlinear $\mathrm{Var}(m_{N})$ of Eq. (6). Here, the $V$-mode inherits the long memory
$\mathrm{Cov}(V_{1},V_{k})=\frac{2H(2H-1)}{1+\rho}\,k^{2H-2}+O(k^{2H-4}),$ (17)
whereas the $U$-mode is two derivatives shorter, at $\mathrm{Cov}(U_{1},U_{k})\propto k^{2H-4}$. Thus, the long-range dependence of $m_{N}-\langle m_{N}\rangle$} is carried entirely by the sum built on the terms $\{\mathrm{He}_{2}(V_{i})\}$,
$m_{N}-\langle m_{N}\rangle\underset{N\rightarrow\infty}{\sim}-\frac{\sqrt{1-\rho^{2}}}{4\pi}\sum_{i=1}^{N-1}\mathrm{He}_{2}(V_{i}).$ (18)
Equation (18) represents, up to negative constant prefactor, the celebrated form of the Rosenblatt random variable [63, 64], where the standard Gaussian variables $V_{i}$ are correlated, $\mathbb{E}(V_{0}V_{k})\propto k^{-D}$ with $0<D<1/2$. Using Eq. (17), we find that D=22HD=2-2H and according to the Dobrushin–Major–Taqqu theorem [61, 63, 64] we finally obtain the key exact result that, for $H>3/4$,
$\frac{m_{N}-\langle m_{N}\rangle}{\sqrt{\mathrm{Var}(m_{N})}}\Longrightarrow-\mathcal{R}_{D},$ (19)
where $\mathcal{R}_{D}$ is the canonical Rosenblatt random variable of unit variance.
For $H\leq 3/4$, Eq. (16) contains only short-range correlated random variables and the Breuer–Major type CLT applies [61, 62]: the fluctuations of $m_{N}$ are Gaussian  [^5] In the marginally non-summable case $H=3/4$ a CLT still holds (but with unusual normalization $\sqrt{N\log N}$), with logarithmically slow convergence to the normal distribution [66]. and the centered and normalized $m_{N}$ converges to the Brownian motion $\mathcal{B}$ [61, 66, 67],
$\frac{m_{N}-\langle m_{N}\rangle}{\sqrt{\mathrm{Var}(m_{N})}}\Longrightarrow\mathcal{B}.$ (20)
Figure 2 showing the one-time PDF of $m_{N}$ obtained from the fBm simulations  [^6] The fBm simulations are performed based on the Davies-Harte method [75, 76, 77], implemented in the fbm package in Python. confirms that $m_{N}$ follow either Gaussian or Rosenblatt distribution and with this the CLT-breaking at $H=3/4$  [^7] We note that computation of the Rosenblatt PDF is an active topic in mathematical literature [81, 82]; here we use the numerically obtained Rosenblatt PDF from Ref. [81]..
Finally, Eqs. (19) and (20) represent the convergence in the process, which provide us the full statistics of $m_{N}$ at any number of times. Here we exemplify it on the case of two times by considering the covariance of
$Z_{K}(t)=\frac{m_{[Kt]}-\langle m_{[Kt]}\rangle}{\sqrt{m_{K}}},$ (21)
which is the basic tool widely used by physicists to describe ageing in the dynamics [70].
For $H>3/4$, $Z_{K}(t)$ converges to the Rosenblatt process $\mathcal{R}(t)$, whose covariance reads [71]
$\mathbb{E}[\mathcal{R}(t)\mathcal{R}(s)]=\frac{1}{2}\big(t^{2(2H-1)}+s^{2(2H-1)}-|t-s|^{2(2H-1)}\big).$ (22)
For $H\leq 3/4$, $\mathbb{E}[Z_{K}(t)Z_{K}(s)]$ converges to that of the standard Brownian motion $\mathbb{E}[\mathcal{B}(t)\mathcal{B}(s)]=\min(t,s)$ [61]. Figure 3 confirms the two distinct behaviors of $\mathbb{E}[Z_{K}(t)Z_{K}(s)]$ depending on $H$ and highlights the fundamental transition of the number of minima in fBm at $H>3/4$.
Acknowledgements. The work was inspired by the talk of G. Schehr in Journées de Physique Statistique 2025 in Paris. We thank the anonymous referee for putting our attention on Refs. [58, 72, 73].

### v4

@@
 $P_{N}(\Phi)$ $=\frac{1}{\sqrt{(2\pi)^{N}\det{\bf R}}}\exp\left(-\frac{1}{2}\Phi^{T}{\bf R}^{-1}\Phi\right),$ (1)
-where $\Phi=(\phi_{0},\dots,\phi_{N-1})$}, and ${\bf R}=(\rho_{ij})$} is the correlation matrix, whose elements depend only on the distnance between increments, $\rho_{ij}=\rho_{0|i-j|}$ and
+where $\Phi=(\phi_{0},\dots,\phi_{N-1})$}, and ${\bf R}=(\rho_{ij})$} is the correlation matrix, whose elements depend only on the distance between increments, $\rho_{ij}=\rho_{0|i-j|}$ and
 $\rho_{0k}=\frac{1}{2}\left(|k-1|^{2H}+|k+1|^{2H}-2k^{2H}\right),$ (2)
@@
 Meanwhile its prefactor is modified, but it depends only on the nearest-neighbor increment correlation $\rho_{01}$.
-In fact, it is the same as for the so-called autoregressive model AR(1) [54, 55], which is a discretized Ornstein-Uhlenbeck process [56], and whose correlation matrix is given by:
+In fact, it is the same as for the so-called autoregressive model AR(1) [60, 61], which is a discretized Ornstein-Uhlenbeck process [62], and for the model with the correlation matrix:
 […]
@@
 $\text{Var}(m_{N})\sim$ $\begin{cases}c_{H}N,&H<\frac{3}{4},\\ \frac{9(\sqrt{2}-1)}{64\pi^{2}}N\log N+b_{3/4}N,&H=\frac{3}{4},\\ \frac{4^{1-H}-1}{2\pi^{2}}\frac{H^{2}(2H-1)}{(4H-3)}N^{4H-2},&H>\frac{3}{4}.\end{cases}$ (6)
-at leading order in $N$ (see Supplemental Material (SM)  [^2] See Supplemental Material at [url] for detailed calculations and numerical checks of asymptotics for detailed calculations and numerical checks of asymptotics, as well as for definitions of coefficients $c_{H}$ and $b_{3/4}$).
+at leading order in $N$ (see Supplemental Material (SM)  [^2] See Supplemental Material at [url] for detailed calculations and numerical checks of asymptotics, which includes Refs.
+[59, 64, 38, 21, 22, 60, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75] for detailed calculations and numerical checks of asymptotics, as well as for definitions of coefficients $c_{H}$ and $b_{3/4}$).
 We note that the asymptotic scaling exponents of Eq.
@@
 Importantly, this behavior comes from the long-range increment correlations $\rho_{0k}$.
-Given that $\rho_{0k}\propto k^{2H-2}$ for $k\rightarrow\infty$ and that the correlation function $\rho_{0k}$ is squared in $\text{Var}(m_{N})\propto\sum_{k=2}^{N-2}(N-k-1)\rho_{0k}^{2}$, the correlations become non-summable when 2(2H2)=12(2H-2)=1, i.e.
+Given that $\rho_{0k}\propto k^{2H-2}$ for $k\rightarrow\infty$ and that the correlation function $\rho_{0k}$ is squared in $\text{Var}(m_{N})\propto\sum_{k=2}^{N-2}(N-k-1)\rho_{0k}^{2}$, the correlations become non-summable when 2(2H2)=12(2H-2)=-1, i.e.
 at $H=3/4$, which is not the usual threshold value $H=1/2$ that separates fBm sub- and superdiffusion  [^3] On the level of a path-integral representation, the fBm action involves a kernel with decay exponent 22H2-2H [77]; here the correlations of the minima involve the squared correlations of fBm increments, which are expected to be controlled by an effective kernel with exponent 44H4-4H (which becomes marginal, equal to $1$, at $H=3/4$)..
@@
 As a result, the crossover to the asymptotic $N\log N$ behavior is extremely slow in practice (see SM).
-The strong correlations of $\{\phi_{i}\}$ leading to the anomalously large (superlinear) fluctuations of $\text{Var}(m_{N})$ have a fundamental consequence for the distribution of $m_{N}$, for which the CLT breaks down.
+The strong correlations of $\{\phi_{i}\}$ leading to the anomalously large (superlinear) fluctuations of $\text{Var}(m_{N})$ have a fundamental consequence for the distribution of $m_{N}$, for which the central limit theorem (CLT) breaks down.
 We show below this breakdown analytically and confirm it numerically through simulations (see Fig.
-2(a))  [^4] The breakdown was also found in the continuous setting of the counterpart zero-crossing problem for fBm studied by Slud, who showed that the CLT fails for $H>3/4$ by representing the number of zero crossings as a second-order Wiener–Itô integral [72, 73]..
+2(a)).
 In the analytic approach, we use the methods Ref.
 [66] for functionals of Gaussian correlated variables, which allow us to go beyond one-time distribution and to show the convergence of $m_{N}$ to the processes belonging to two different universality classes (Rosenblatt and Brownian motion), separated through the threshold value $H=3/4$.
+Note that, in the continuous setting, as outlined in the review by Kratz (Th.2.27) [78], the limiting process of the related problem of zero crossings’ number is given in the supercorrelated regime through a form involving implicitly the correlation function and is expressed through the multiple Wiener integral of the second order.
+This fundamental mathematical object, originally found by Slud [69], represents the Rosenblatt process, but practically is of little use.
+Here, in the discrete setting, we construct the Rosenblatt process in closed form with explicit analytic parametrisation.
 First step consists in the expansion of $m_{N}$ in the probabilists’ Hermite polynomials $\mathrm{He}_{n}(\phi_{i})$. Using
@@
 Figure 2: (a) Probability density function (PDF) of centered and normalized $m_{N}$ for N=1024{N=1024}, obtained from simulations (symbols), compared with Gaussian PDF (black line) and (b) for $N=10^{6}$} and various $H>3/4$ compared with the PDF of the negative Rosenblatt variable (lines).
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 3: The covariance of the number of minima $\mathbb{E}[Z_{K}(t)Z_{K}(s)]$ (symbols) compared with theory (solid lines), For (a)-(c) the theoretical covariance is $\min(t,s)$ of the Brownian motion and for (d)-(f) is given by Eq.
+(22) for the Rosenblatt process.
 Because the rank $2$ terms drive the CLT breaking, we are going now to analyse Eq.
-(9), by spanning the quadratic part $M_{i}^{\circ}\equiv M_{i}-\langle M_{i}\rangle$ on the orthogonal vectors of the complete quadratic Wick basis
+(9), by spanning the quadratic part $M_{i}^{\circ}\equiv M_{i}-\langle M_{i}\rangle$ on the linearly independent vectors of the complete quadratic Wick basis
 $B_{1}\equiv\mathrm{He}_{2}(\phi_{i-1})+\mathrm{He}_{2}(\phi_{i})=\phi_{i-1}^{2}+\phi_{i}^{2}-2,$
@@
 $Q_{i}=\frac{\rho(\phi_{i-1}^{2}+\phi_{i}^{2}-2)-2(\phi_{i-1}\phi_{i}-\rho)}{4\pi\sqrt{1-\rho^{2}}}.$ (13)
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Figure 3: The covariance of the number of minima $\mathbb{E}[Z_{K}(t)Z_{K}(s)]$ (symbols) compared with theory (solid lines), For (a)-(c) the theoretical covariance is $\min(t,s)$ of the Brownian motion and for (d)-(f) is given by Eq.
-(19) for the Rosenblatt process.
 Introducing the normalized sum and difference modes
@@
 Here, the $V$-mode inherits the long memory
-$\mathrm{Cov}(V_{1},V_{k})=\frac{2H(2H-1)}{1+\rho}\,k^{2H-2}+O(k^{2H-4}),$ (17)
+$\mathrm{Cov}(V_{1},V_{k+1})=\frac{2H(2H-1)}{1+\rho}\,k^{2H-2}+O(k^{2H-4}),$ (17)
 whereas the $U$-mode is two derivatives shorter, at $\mathrm{Cov}(U_{1},U_{k})\propto k^{2H-4}$.
@@
 Here we exemplify it on the case of two times by considering the covariance of
-$Z_{K}(t)=\frac{m_{[Kt]}-\langle m_{[Kt]}\rangle}{\sqrt{m_{K}}},$ (21)
+$Z_{K}(t)=\frac{m_{[Kt]}-\langle m_{[Kt]}\rangle}{\sqrt{\mathrm{Var}(m_{K})}},$ (21)
 which is the basic tool widely used by physicists to describe ageing in the dynamics [82].
@@
 Acknowledgements. The work was inspired by the talk of G. Schehr in Journées de Physique Statistique 2025 in Paris. We thank the anonymous referee for putting our attention on Refs. [65, 69, 78].
+Data availability. The data that support the findings of this article are openly available [83].